OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 5x − 3y − z if x = −2, y = 2, and z = −3 −19 19 13 −13

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

here u just have to substitute the values for each variable, like this: 5x − 3y − z if x = −2, y = 2, and z = −3 5(-2)-3(2)-(-3) = -10-6+3 = -13

OpenStudy (anonymous):

5(-2)-3(2)-(-3) -10-6+3 -16+3=-13<<<<<<answer